# Spin correlations in associated top-quark and dark matter production at Large Hadron Collider

### v1

@@ Abstract
In the paper, we consider the processes of single top quark production in association with dark matter particles through the t-channel in simplified models with scalar and pseudoscalar mediators. Within the framework of these models, we analyze the spin correlations that arise in the top quark production process in association with dark matter, comparing them with the production of a top quark in the Standard Model. We consider angular correlations of kinematic variables, which define admissible regions in momentum space and can be used experimentally to distinguish between the fraction of the total momentum carried away by the scalar or pseudoscalar mediator.
@@
Dark matter; top-quark; spin correlations.
@@ 1 Introduction
Convincing astrophysical and cosmological observations indicate the presence of new, weakly interacting, neutral substance that called Dark Matter (DM). Under the assumption that dark matter has the nature of particles, various extensions of the Standard Model (SM) are created to explain the origin of DM and its possible detection in an experiment.
@@
The presence of non-SM particles can change the known distributions of kinematic variables in the SM. In this work, we analyze spin correlations in the single top quark production in association with a scalar and pseudoscalar DM mediators. It would seem obvious that the kinematic distributions associated with spin correlations should not change, at least in the case of a scalar mediator. However, the presented analysis shows that this is not the case. An example of a kinematic variable associated with spin correlations, which is affected by a mediator, including a scalar one, is given. A number of proposed kinematic variables are used to train a neural network that separates the DM mediator and SM neutrino contributions to the total missing energy determined at the detector. It is shown that multivariate analysis techniques such as deep neural networks can help improve the sensitivity of the analysis to the presence of DM mediators. In the framework of the considered simplified models with various mediators, in this article we focused on the spin properties of the t-channel single top quark production in association with the DM mediators.
The paper is organized as follows. Section 2 describes two classes of phenomenological models that predict the processes of top-quark production in association with DM particles and gives an overview of the latest results of the search for DM particles in association with top-quarks. The section 3 describes and summarizes spin correlations in the processes of associative production of dark matter with top quarks at the LHC. In Section 4 the analytical expressions for mediator momentum components are provided and additional methods of improving the manifestation of spin correlations are discussed. The section 5 gives the final results. The conclusion describes possible applications of the proposed method for increasing the sensitivity for DM searches at the LHC.
@@
3 Spin correlations in associated production of top-quark and dark matter at LHC
This article focuses on single-top quark production with subsequent leptonic decay, which is well studied in the SM. One of the significant properties of such process in SM is correlation of the spin of top-quark and its decay products with the direction of the down-type quark. From that one can get the relation for the cosine of angle between down-type quark and the lepton in a specific case of $u\bar{d}\rightarrow t\bar{b}$ and a subsequent $t\rightarrow\bar{l}\nu b$ decay [21].
@@ 3 Spin correlations in associated production of top-quark and dark matter at LHC
Since the mediator is a scalar particle, changes to this rule were not expected. To check this property events with DM mediator were generated in CompHEP using a scalar mediator model. However, as can be seen on Fig. 1, addition of the DM mediator breaks the spin correlations. This is caused due to the cosine being constructed in the top-quark rest frame (TRF), which is changed with the additional emission of a heavy scalar boson. Since all particles are available at a parton level generation, we have added the mediator to the reconstruction of the missing transverse energy, which restored the expected distribution.
Figure 1: Distribution of cosine of angle between lepton and down-type quark in top-quark rest frame in processes $pp\rightarrow t(\rightarrow\nu_{l}\bar{l}b)q\ (SM)$ and $pp\rightarrow t(\rightarrow\nu_{l}\bar{l}b)q\bar{\chi}\chi$ (neutrino TRF) with presence of a scalar mediator in a model with $m_{\chi}=1\ GeV,\ m_{\phi}=400\ GeV,\ g_{\chi}=g_{\nu}=1$. For comparison the same distribution in a top-quark emitting the DM mediator is also provided (neutrino and mediator TRF).
@@
Next, all events were hadronized in Pythia8 [22] and run through a detector response simulation with Delphes [23]. As can be seen in Fig. 3, distributions still clearly differ, which allows an additional observable for the separation of DM events in single top-quark production based on spin correlations.
Figure 2: Distribution of cosine of angle between lepton and down-type quark in top-quark rest frame in processes $pp\rightarrow t(\rightarrow\nu_{l}\bar{l}b)q\ (SM)$ and $pp\rightarrow t(\rightarrow\nu_{l}\bar{l}b)q\bar{\chi}\chi$(neutrino TRF) with presence of a pseudoscalar mediator in a model with $m_{\chi}=1\ GeV,\ m_{\phi}=400\ GeV,\ g_{\chi}=g_{\nu}=1$. For comparison the same distribution in a top-quark emitting the DM mediator is also provided (neutrino and mediator TRF)).
@@ 4.2 Neutrino momentum reconstruction with neural networks
Another opportunity to utilize spin correlations lies in learning the reconstruction of neutrino and DM mediator momenta from the MC simulation directly. This task can be solved with the application of various machine learning methods. For this work, we have trained a Multi-Layer Perceptron in a multi-output regression setting for two separate test cases:
@@
We tune the following parameters of the training pipeline: the width, the depth of the network, hidden layer activations, learning rate, weight decay and dropout probability. We also jointly tune the preproccessing pipeline, varying the scaler (No scaler, MinMax, Standard, Power, Quantile) of the input and target variables independently. In order to maintain a fair comparison, all output variables are inversely scaled after the inference, and are compared with the original. We use Mean Absolute Percentage Error as the tuning criterion. The resulting network is trained until the validation loss stops decreasing.
@@ 4.2.1 SM network
For the SM network only the SM sample from CompHEP is used. This is intended as a proof-of-concept case for our method of momentum reconstruction. As can be seen from Fig. 5 and 6, reconstructed distributions of both the components of the neutrino momentum and $cos\theta_{\bar{d}\bar{\ell}}$ perfectly align with the desired values obtained from MC.
@@ 5 Results
In this work we have presented a promising new observable to distinguish events containing scalar of pseudoscalar mediator of dark matter based on the spin correlations of the top-quark in SM. The new angular variable discussed in Section 3 provides a significant separation between SM and DM processes. In combination with separation by missing transverse momentum this variable can be used to improve search for dark matter at various colliders. In order to further enhance the separation power of the method two approaches were tested: analytical approach based on the kinematic relations shown in Section 4.1 and the machine learning approach using neural networks for the reconstruction of 4-momenta described in Section 4.2. Both approaches did not achieve desired improvements; however, the approach using neural networks shows significant potential and, probably, can be improved with more sophisticated NN architectures.

### v2

# Angular correlations in associated top-quark and dark matter production at Large Hadron Collider

@@ Abstract
In the paper, we consider the processes of single top-quark production in association with dark matter particles through the t-channel in simplified models with scalar and pseudoscalar mediators. Within the framework of these models, we analyze the angular correlations that arise in the top quark production process in association with dark matter, comparing them with the production of a top quark in the Standard Model. We consider angular correlations of kinematic variables, which define admissible regions in momentum space and can be used experimentally to distinguish between the fraction of the total momentum carried away by the scalar or pseudoscalar mediator. We demonstrate that scalar mediator significantly changes the rest frame of the well-known angular correlation observable and changes the distributions. In this case, this variable provides clear separation between the SM and DM contribution and can be used to increase the sensitivity of future DM analyses.
@@
Dark matter; top-quark; angular correlations.
@@ 1 Introduction
Convincing astrophysical and cosmological observations indicate the presence of new, weakly interacting, neutral substance called Dark Matter (DM). Under the assumption that dark matter has the nature of particles, various extensions of the Standard Model (SM) are created to explain the origin of DM and its possible detection in an experiment.
@@
The presence of non-SM particles can change the known distributions of kinematic variables in the SM. In single top production, the top quark is highly polarized due to (V-A) electroweak vertices in production and decay of the top quark. The $P_{z}$ component of the polarization vector is proportional to the cosine of the angle between the spectator quark and the charged lepton in the top-quark decay chain, in the rest frame of the top-quark [19]. This angular correlation is a well-known property and is very sensitive to deviations in Wtb interactions. The production of scalar dark matter mediator from the top-quark-line does not change the polarization of the top-quark, but the direction of the d-quark momentum is not now the axis along which the top quark is about 100% polarized, and as a result, changes the observable cosine.
In this work, we analyze angular correlations in the single top quark production in association with a scalar and pseudoscalar DM mediators. It would seem obvious that the kinematic distributions associated with angular correlations should not change, at least in the case of a scalar mediator. However, the presented analysis shows that this is not the case. An example of a kinematic variable associated with angular correlations, which is affected by a mediator, including a scalar one, is given. A number of proposed kinematic variables are used to train a neural network that separates the DM mediator and SM neutrino contributions to the total missing energy determined at the detector. It is shown that multivariate analysis techniques such as deep neural networks can help improve the sensitivity of the analysis to the presence of DM mediators. In the framework of the considered simplified models with various mediators, in this article we focused on the spin properties of the t-channel single top quark production in association with the DM mediators.
The paper is organized as follows. Section 2 describes two classes of phenomenological models that predict the processes of top-quark production in association with DM particles and gives an overview of the latest results of the search for DM particles in association with top-quarks. The section 3 describes and summarizes angular correlations in the processes of associative production of dark matter with top quarks at the LHC. In Section 4 the analytical expressions for mediator momentum components are provided and additional methods of improving the manifestation of angular correlations are discussed. The section 5 gives the final results. The conclusion describes possible applications of the proposed method for increasing the sensitivity for DM searches at the LHC.
@@
Typical Feynman diagram of the studied process is presented in the Figure 1.
Figure 1: Representative Feynman diagram of the $q^{\prime}g\rightarrow t\bar{b}q$ process with the presence of scalar/pseudoscalar mediator.
3 Angular correlations in associated production of top-quark and dark matter at LHC
This article focuses on single-top quark production with subsequent leptonic decay, which is well studied in the SM. One of the significant properties of such process in SM is correlation of the spin of top-quark and its decay products with the direction of the down-type quark. From that one can get the relation for the cosine of angle between down-type quark and the lepton in a specific case of $ug\rightarrow t\bar{b}d$ and a subsequent $t\rightarrow\bar{l}\nu b$ decay [19]. Since the mediator is a scalar particle, changes to this rule were not expected. For the top-quark rest frame (TRF), firstly, 4-momentum of the top-quark is calculated: as a sum of 4-momenta of neutrino, charged lepton and the b-quark for the “neutrino” TRF case and with the addition of mediator in the “neutrino and mediator” TRF case. Secondly, charged lepton and down-type quark are boosted to the TRF using the previously discussed top-quark momentum. Then, the cosine of the angle between the charged lepton and down-type quark is calculated as the normalized dot product of corresponding momenta. Thus, the distribution changes due to the inclusion of the heavy mediator in the TRF calculations in the first step.
@@ 3 Angular correlations in associated production of top-quark and dark matter at LHC
To check this property events with DM mediator were generated in CompHEP using a scalar mediator model. However, as can be seen on Fig. 2, addition of the DM mediator breaks the angular correlations. This is caused due to the cosine being constructed in the top-quark rest frame (TRF), which is changed with the additional emission of a heavy scalar boson. Since all particles are available at a parton level generation, we have added the mediator to the reconstruction of the missing transverse energy, which restored the expected distribution.
Figure 2: Distribution of cosine of angle between lepton and down-type quark in top-quark rest frame in processes $pp\rightarrow t(\rightarrow\nu_{l}\bar{l}b)q\ (SM)$ and $pp\rightarrow t(\rightarrow\nu_{l}\bar{l}b)q\bar{\chi}\chi$ (“neutrino” TRF) with presence of a scalar mediator in a model with $m_{\chi}=1\ GeV$, 2 samples with $m_{\phi}=400\ GeV$ and $1500\ GeV,\ g_{\chi}=g_{\nu}=$ 1. For comparison the same distribution in a top-quark emitting the DM mediator is also provided (“neutrino and mediator” TRF).
@@
Next, all events were hadronized in Pythia8 [22] and run through a detector response simulation with Delphes [23]. The selection criteria for Delphes are: 1 lepton with $P_{t}>26$ GeV, $\eta<2.4$, 2 or 3 jets with $P_{t}>30$ GeV, exactly one of which is b-tagged, $MET>20$ GeV. As can be seen in Fig. 4, distributions still clearly differ, which allows an additional observable for the separation of DM events in single top-quark production based on angular correlations.
Figure 3: Distribution of cosine of angle between lepton and down-type quark in top-quark rest frame in processes $pp\rightarrow t(\rightarrow\nu_{l}\bar{l}b)q\ (SM)$ and $pp\rightarrow t(\rightarrow\nu_{l}\bar{l}b)q\bar{\chi}\chi$(neutrino TRF) with presence of a pseudoscalar mediator in a model with $m_{\chi}=1\ GeV$, 2 samples with $m_{\phi}=400\ GeV$ and $1500\ GeV,\ g_{\chi}=g_{\nu}=$ 1. For comparison the same distribution in a top-quark emitting the DM mediator is also provided (neutrino and mediator TRF)).
@@ 4.2 Neutrino momentum reconstruction with neural networks
Another opportunity to utilize angular correlations lies in learning the reconstruction of neutrino and DM mediator momenta from the MC simulation directly. This task can be solved with the application of various machine learning methods. For this work, we have trained a Multi-Layer Perceptron in a multi-output regression setting for two separate test cases:
@@
The Standard Scaler is used for all input and output variables; in order to maintain a fair comparison, all output variables are inversely scaled after the inference, and are compared with the original. We use Mean Absolute Percentage Error as the tuning criterion. The resulting network is trained until the validation loss stops decreasing.
@@ 4.2.1 SM network
For the SM network only the SM sample from CompHEP is used. This is intended as a proof-of-concept case for our method of momentum reconstruction. As can be seen from Fig. 6 and 7, reconstructed distributions of both the components of the neutrino momentum and $cos\theta_{\bar{d}\bar{\ell}}$ perfectly align with the desired values obtained from MC.
@@ 5 Results
In this work we have presented a promising new observable to distinguish events containing scalar of pseudoscalar mediator of dark matter based on the angular correlations of the top-quark in SM. The new angular variable discussed in Section 3 provides a significant separation between SM and DM processes. In combination with separation by missing transverse momentum this variable can be used to improve search for dark matter at various colliders. In order to further enhance the separation power of the method two approaches were tested: analytical approach based on the kinematic relations shown in Section 4.1 and the machine learning approach using neural networks for the reconstruction of 4-momenta described in Section 4.2. Both approaches did not achieve desired improvements; however, the approach using neural networks shows significant potential and, probably, can be improved with more sophisticated NN architectures.